OpenStudy (anonymous):

5. A sound wave is modeled with the equation y=1/4 cos 2pi/3 θ. a. Find the period. Explain your method. (I think its 3) b. Find the amplitude. Explain your method. (I think its1/4) c. What is the equation of the midline? What does it represent?

OpenStudy (anonymous):

@Teddyiswatshecallsme @fallenangelorchid

OpenStudy (anonymous):

@fani1996

OpenStudy (fallenangelorchid):

Sorry, I dunno.

OpenStudy (anonymous):

im sorry to @fallenangelorchid i dunno either lol

OpenStudy (anonymous):

@danish071996

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

@uri @undeadknight26

OpenStudy (anonymous):

@sammixboo

OpenStudy (shamim):

The standard equation of a sound wave is y=A*sin(omega*t)

OpenStudy (shamim):

I think ur given equation is y=(1/4)*sin(2*pi/3)t

OpenStudy (shamim):

Right???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry it took me so long to answer os is laggy for me

OpenStudy (shamim):

U wrote theta in ur given equation. I think that was wrong!!!

OpenStudy (shamim):

Here A=amplitude of wave

OpenStudy (shamim):

Snyway if u compare ur equation with my standard equation u will get A=1/4 Right?

OpenStudy (shamim):

Omega =2*pi/3 2*pi/T=2*pi/3 T=?

OpenStudy (anonymous):

so its 3pie/2 right

OpenStudy (anonymous):

theta was in my question man

OpenStudy (shamim):

No

OpenStudy (shamim):

T=3 Right?

OpenStudy (anonymous):

idk i have nvr done this before my teacher just dropped it on me and didnt explain what to do

OpenStudy (shamim):

It was (2*pi)/3=(2*pi)/T T=3 Right?

OpenStudy (shamim):

Do u need more explaination?

OpenStudy (anonymous):

plz